OpenStudy (anonymous):

can you give me an example of 4 equation and 4 unknown. i need step by step please.. i really need this now.. seriously.

OpenStudy (anonymous):

w+2x-2y+3z=12 2w+3x+y+2z=10 3w-x+3y+4z=16 -w+4x-2y+4z=20

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{green}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{purple}{Welcome~to~\color{purple}{Open}}\color{purple}{Study!!!!}\) Are those your equations that you need to solve?

OpenStudy (anonymous):

thank you and yes thats the equation i need to solve..

OpenStudy (yanasidlinskiy):

I will NOT give you the answers. We'll work some out together and then you'll do the rest on your own. And then, I'll check it for you:) And your Welcome!!:) So, lets start with the first equation. What's the first step do we need to do?

OpenStudy (anonymous):

can u just give me some examples? of 4 equation and 4 unknown?

OpenStudy (yanasidlinskiy):

You provided me with 4 equations and 4 unknowns yourself. Now we just have to solve them. Or, do we have to solve them?

OpenStudy (anonymous):

i really dont know if that equation is 4 equation and 4 unknown my teacher told me to solve that equation and someone told me its 4 equation and 4 unknown.. so here i am asking u guys can u please give me some examples of 4 equation and 4 unknown . and ill try solve my equation all by myself~

OpenStudy (alekos):

this is best done using the augmented matrix method

OpenStudy (anonymous):

were not arguing though but thanks.